# Classification of abelian Schur groups I

### v1

@@ 7.2. Nondense $S$-rings
    $4\nmid|G/L|$ and $|G/U|\neq 2$.
The next proposition provides a refinement of one of the cases from Proposition 7.3.
Proposition 7.5.
Suppose that $H$ is not an $\mathcal{A}$-subgroup and $\mathcal{A}$ is a nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ with $4\nmid|U|$ and $|L|\neq 2$. Then $\mathcal{A}$ is a nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}$-section $U_{1}/L_{1}$ satisfying one of the following conditions:
  1. (1)
    $|U_{1}/L_{1}|\leq 2$;
  2. (2)
    $\mathcal{A}_{L_{1}}$ is $\otimes$-complemented in $\mathcal{A}_{U_{1}}$;
  3. (3)
    $U_{1}\leq U$ and $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$.
@@
Note that $U$ is cyclic because $4\nmid|U|$. If $|\operatorname{rad}(\mathcal{A}_{U})|=1$, then $U_{1}/L_{1}=U/L$ satisfies Condition $(3)$ of Proposition 7.5 and we are done. Further, we assume that
$|\operatorname{rad}(\mathcal{A}_{U})|>1.$
Suppose that $|U|$ is odd. Then $|U|=pq$ or $|U|$ is a $p$-power. In the former case, $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=L/L$-wreath product and $|L|$ is prime by Lemma 6.12. So $\mathcal{A}$ is the $U_{1}/L_{1}=L/L$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(1)$ from Proposition 7.5. In the latter one, $\mathcal{A}_{U}$ is the nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}_{U}$-section $U_{1}/L_{1}$ of $U$ such that $L_{1}$ is the least nontrivial $\mathcal{A}_{U}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$ by Lemma 6.10. By the definitions of $U_{1}$ and $L_{1}$, we have $U_{1}\leq U$ and $L_{1}\leq L$. This implies that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(3)$ as required.
Now suppose that $|U|$ is even. Then $|U|=2m$, where $m$ is odd. Let $A$ and $W$ be the Sylow $2$-subgroup and the Hall $2^{\prime}$-subgroup of $U$, respectively. If $|W|=1$, then $|L|=|U|=2$ and Condition $(1)$ from Proposition 7.5 holds for $U_{1}/L_{1}=U/L$. Further, we assume that $|W|>1$. The least $\mathcal{A}_{U}$-subgroup containing $A$ and a maximal $\mathcal{A}_{U}$-subgroup of $W$ are denoted by $A_{1}$ and $W_{1}$, respectively. If $|W_{1}|=1$, then $\mathcal{A}_{U}$ is trivial or the $A_{1}/A_{1}$-wreath product by Lemma 5.3 applied to $\mathcal{A}_{U}$, $W_{1}$, and $A_{1}$. In the former case, $U=L$ and $U_{1}/L_{1}=U/L$ satisfies Condition $(1)$. In the latter one, $A_{1}$ is the least $\mathcal{A}_{U}$-subgroup because otherwise $|W_{1}|>1$ and hence $A_{1}\leq L$. So $\mathcal{A}$ is the $U_{1}/L_{1}=A_{1}/A_{1}$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(1)$.
Further, we assume that
$|W_{1}|>1.$
Observe that $M\leq L$ for some minimal $\mathcal{A}_{W_{1}}$-subgroup $M$ because $|L|\neq 2$. One can see that $\mathcal{A}_{W_{1}}$ is the $V/M$-wreath product (possibly, trivial), where $V$ is an $\mathcal{A}_{W_{1}}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{V})|=1$. Indeed, this follows from Lemma 6.2 and Lemma 6.12 if $|W_{1}|=pq$ and from Lemma 6.2 and Lemma 6.10 if $|W_{1}|$ is a $p$-power. Lemma 5.4 implies that
Let $W_{1}A_{1}=U$. If $W_{1}\cap A_{1}$ is trivial, then $\mathcal{A}_{A_{1}}$ is primitive and hence $\mathcal{A}_{A_{1}}=\mathcal{T}_{A_{1}}$ by Lemma 6.1. Eq. (13) implies that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\otimes\mathcal{A}_{A_{1}}$ and consequently $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=(V\times A_{1})/M$-wreath product (possibly, trivial). As $M\leq L$, we conclude that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and
we obtain $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$. Thus, $U_{1}/L_{1}$ satisfies Condition $(3)$ and we are done.
Further, we assume that $W_{1}A_{1}<U$. This implies that
$W_{1}<W.$
In this case, $|W_{1}|$ is a prime power and hence $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup. Lemma 5.3 yields that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\wr\mathcal{A}_{U/W_{1}}$ or $\mathcal{A}_{U}=\mathcal{A}_{W_{1}A_{1}}\wr_{(W_{1}A_{1})/A_{1}}\mathcal{A}_{U/A_{1}}$. In the former case, $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=V/M$-wreath product because $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$ and $M\leq L$. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$, the section $U_{1}/L_{1}$ satisfies Condition $(3)$ and we are done.
Suppose that
If $\mathcal{A}_{A_{1}}\neq\mathcal{T}_{A_{1}}$, then $|A_{1}\cap W_{1}|>1$. Due to Eqs. (13) and (14), we conclude that $\mathcal{A}_{U}$ is the nontrivial $W_{1}/(W_{1}\cap A_{1})$-wreath product. Since $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup, we have $M\leq A_{1}\cap W_{1}$ and consequently $\mathcal{A}_{U}$ is the $W_{1}/M$-wreath product. As $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$, we conclude that $\mathcal{A}_{U}$ is the nontrivial $V/M$-wreath product. Observe that $\mathcal{A}$ is the $U_{1}/L_{1}=V/M$-wreath product because $M\leq L$. The section $U_{1}/L_{1}$ satisfies Condition $(3)$ because $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and we are done.
(15) $\mathcal{A}_{W_{1}A_{1}}=\mathcal{A}_{W_{1}}\otimes\mathcal{T}_{A_{1}}.$
If $A_{1}\neq A$, then $|W_{1}A_{1}|=2pq$, $U=W_{1}A_{1}$, and $L\in\{W_{1},A_{1}\}$. So $U_{1}/L_{1}=U/L$ satisfies Condition $(2)$ by Eq. (15) and we are done. Further, we assume that $A_{1}=A$.
@@
With the above notation, let $X\in\mathcal{S}(\mathcal{A})$ such that $X_{D}\nsubseteq W_{1}$. Then $A\leq\operatorname{rad}(X)$.
@@
If $X\subseteq U$, then we are done by Eq. (14). Further, we assume that $X\nsubseteq U$ and hence $L\leq R_{0}=\operatorname{rad}(X)$. Let $R=\langle X\rangle$. One can see that
$R\cap(W\setminus W_{1})\neq\varnothing$
because $X_{D}\nsubseteq W_{1}$. Together with Eq. (14), this implies that $R\geq A$. The image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/R_{0}$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. Clearly, $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1$ and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable.
Assume that $A\nleq R_{0}$. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu\geq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Due to Eq. (14), we have $R_{0}\cap D\leq W_{1}$ and
for every $Y\in\mathcal{S}(\mathcal{A})_{R}$ inside $R\cap(U\setminus(W_{1}A))$. Observe that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial because $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$ is a proper nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup. If $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is cyclic, then $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is cyclotomic or a nontrivial tensor product by Lemma 6.2. However, this contradicts to Eq. (16). Suppose that $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is noncyclic. Then in view of Eq. (16), the Hall $2^{\prime}$-subgroup of $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is not an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup and hence Proposition 7.4 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$. Since $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial and indecomposable, we conclude that
for some $\mathcal{A}_{R}$-subgroups $R_{1}$ and $R_{2}$. As $R_{0}\geq L$ and $|L|\neq 2$, we may assume without loss of generality that $|\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}|$ is a power of $2$. Due to Eqs. (16) and (17), $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}\ngeq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Therefore $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}$ is a cyclic group containing $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Lemma 6.2 and Eq. (16) yield that $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}})|\neq 1$. Together with Eq. (17), this implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is decomposable, a contradiction. ∎
From Lemma 7.6 it follows that $\mathcal{A}$ is the $(W_{1}A)/A$- or $(W_{1}H)/A$-wreath product. In the first case, $\mathcal{A}_{A}$ is $\otimes$-complemented in $\mathcal{A}_{W_{1}}$ by Eq. (15). So $U_{1}/L_{1}=(AW_{1})/A$ satisfies Condition $(2)$ and we are done.
In the second case, $HW_{1}$ is an $\mathcal{A}$-subgroup. Let $a$ be a nontrivial element of $A$, $b\in H\setminus A$, and $T$ a basic set of $\mathcal{A}$ containing $b$. The $S$-ring $\mathcal{A}_{W_{1}H}$ satisfies [13, Hypothesis $5.1$]. From [13, Lemma $5.5$] and the inclusion $\{a\}=A^{\#}\in\mathcal{S}(\mathcal{A})$ it follows that
$T=W_{0}b~\text{or}~T=W_{0}Ab$
for some $\mathcal{A}_{W_{1}}$-subgroup $W_{0}$. As $H$ is not an $\mathcal{A}$-subgroup, $|W_{0}|>1$. If $T=W_{0}Ab$, then $\mathcal{A}_{W_{1}H}$ is the $(W_{1}A)/(W_{0}A)$-wreath product by [13, Lemma $5.4$] and hence $\mathcal{A}$ is the $(W_{1}A)/A$-wreath product and we are done by the previous paragraph. If $T=W_{0}b$, then $W_{2}=\langle T,W_{1}\rangle$ is an $\mathcal{A}_{W_{1}H}$-subgroup such that $|W_{2}\cap A|=1$ and $W_{1}H=W_{2}\times A$. Lemma 3.2(2) implies that $\mathcal{A}_{W_{1}H}=\mathcal{A}_{W_{2}}\otimes\mathcal{A}_{A}$. Therefore $U_{1}/L_{1}=(W_{1}H)/A$ satisfies Condition $(2)$ and we are done. ∎
The case when $\mathcal{A}$ is a nontrivial $U/L$-wreath product satisfying Condition $(3)$ from Proposition 7.4 is dual to the case from Proposition 7.5. So the next proposition is a corollary of Proposition 7.5 and Lemma 4.1.
Proposition 7.7.
Suppose that $H$ is not an $\mathcal{A}$-subgroup and $\mathcal{A}$ is a nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ with $4\nmid|G/L|$ and $|G/U|\neq 2$. Then $\mathcal{A}$ is a nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}$-section $U_{1}/L_{1}$ satisfying one of the following conditions:
  1. (1)
    $|U_{1}/L_{1}|\leq 2$;
  2. (2)
    $\mathcal{A}_{U_{1}/L_{1}}$ is $\otimes$-complemented in $\mathcal{A}_{G/L_{1}}$
  3. (3)
    $L_{1}\geq L$ and $|\operatorname{rad}(\mathcal{A}_{G/L_{1}})|=1$.
7.3. Dense $S$-rings
The lemma below easily follows from computer calculations using [14].
Lemma 7.8.
Every $S$-ring over $H\cong E_{4}$ is normal and can be presented in the form $\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order at most $3$.
Remark 7.9.
With the notation of Lemma 7.8, it is easy to see that $\operatorname{Cyc}(K_{H},H)=\mathbb{Z}H$ if $|K_{H}|=1$, $\operatorname{Cyc}(K_{H},H)\cong\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2}$ if $|K_{H}|=2$, and $\operatorname{Cyc}(K_{H},H)=\mathcal{T}_{H}$ if $|K_{H}|=3$.
Further throughout this subsection, we assume that $H$ and $D$ are $\mathcal{A}$-subgroups. Denote two distinct elements from $H$ by $a$ and $b$. Given $X\in\mathcal{S}(\mathcal{A})$, put $\lambda_{X}=|X\cap Hx|$, where $x\in X$. Due to Lemma 2.1, $\lambda_{X}$ does not depend on $x$.
Lemma 7.10.
Let $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Then $\lambda_{X}=1$.
Proof.
@@ 7.3. Dense $S$-rings
Clearly, $\lambda_{X}\leq|X_{H}|$ and the equality is attained if and only if $X=X_{H}\times X_{D}$. Therefore the lemma holds if $|X_{H}|\leq 2$. So in view of Lemma 3.2(1) and Remark 7.9, we may assume that $X_{H}=H^{\#}=\{a,b,ab\}$. To finish the proof, it remains to show that $\lambda_{X}\neq 2$. Assume the contrary. Then the set $Y=(X_{H}\times X_{D})\setminus X$ is an $\mathcal{A}$-set containing exactly one element from $Hx$ for every $x\in X_{D}$. Together with Lemma 3.2(1) and the equality $X_{H}=H^{\#}$, this implies that $Y$ is exactly a basic set of $\mathcal{A}$ with
@@
Let $x,x^{\prime}\in X_{D}$ of the same order. Then there is a positive integer $m$ coprime to $2n$ such that $x^{n}=x^{\prime}$. Since $\lambda_{X}=2$, the sets $H_{x}=Xx^{-1}\cap H$ and $H_{x^{\prime}}=X(x^{\prime})^{-1}\cap H$ are subsets of $H^{\#}$ of size $2$. Therefore there is $h\in H_{x}\cap H_{x^{\prime}}$. One can see that $(hx)^{m}=hx^{\prime}\in X^{(m)}\cap X$. So $X^{(m)}=X$ by Lemma 2.3. This yields that
@@
Observe that $|Y_{h_{0}}|=|Y_{h}|$ by Lemma 2.1. In view of Eqs. (19) and (20), to obtain a contradiction, it is enough to show that
@@
The set $X_{D}$ is a basic set of $\mathcal{A}_{D}$ (Lemma 3.2(1)). If $|X_{D}\setminus D_{l}^{*}|=0$, then Eq. (21) is obvious. Suppose that $|X_{D}\setminus D_{l}^{*}|>0$, i.e. $X_{D}$ is nonregular. If $n=pq$, then Lemma 6.12 implies that $\mathcal{A}_{D}$ is trivial or a nontrivial generalized wreath product. So $X_{D}=D^{\#}$ or $X_{D}$ is a union of some $L$-cosets, where $L=\operatorname{rad}(X_{D})$ is a nontrivial $\mathcal{A}_{D}$-subgroup. In the former case, $|X_{D}\cap D_{l}^{*}|=(p-1)(q-1)$ and $|X_{D}\setminus D_{l}^{*}|=p+q-2$, whereas in the latter one $|X_{D}\cap D_{l}^{*}|=r(|L|-1)$ and $|X_{D}\setminus D_{l}^{*}|=r$, where $r=|X_{D}/L|$. If $n=p^{k}$, then due to Lemma 6.11, we obtain $X_{D}=U\setminus L$ for some $\mathcal{A}$-subgroups $U\cong C_{p^{l}}>L$ and consequently $|X_{D}\cap D_{l}^{*}|=p^{l-1}(p-1)$ and $|X_{D}\setminus D_{l}^{*}|=p^{l-1}-|L|$. Since $p$ and $q$ are odd, Eq. (21) holds in all cases and we are done. ∎
Proposition 7.11.
Suppose that $|\operatorname{rad}(\mathcal{A}_{D})|=1$ and $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$. Then the following statements hold:
@@
If $\mathcal{A}_{H}=\mathbb{Z}H$, then $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ by Lemma 3.2(2), a contradiction to the assumption of the proposition. So we may assume that $\mathcal{A}_{H}\neq\mathbb{Z}H$. Then Lemma 7.8 and Remark 7.9 imply that $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order $2$ or $3$, and a unique nonsingleton basic set of $\mathcal{A}_{H}$ is of size $|K_{H}|$. We divide the rest of the proof into two cases depending on whether $\mathcal{A}_{D}$ is cyclotomic or not.
Case 1: $\mathcal{A}_{D}$ is noncyclotomic. In view of Lemma 6.2, the $S$-ring $\mathcal{A}_{D}$ is trivial and hence $D^{\#}\in\mathcal{S}(\mathcal{A}_{D})$. Moreover, $n$ is not prime by Lemma 6.7. So further, we assume that $n=pq$ or $n=p^{k}$ for $k\geq 2$. If $n=pq$, then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$ and $D$ is the least $\mathcal{A}$-subgroup whose order is divisible by $p$. By Lemma 5.3 applied to $\mathcal{A}$, $H$, and $D$, we have $\mathcal{A}=\mathcal{A}_{H}\star\mathcal{A}_{D}$. Since $H\cap D$ is trivial, we obtain $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, a contradiction to the assumption of the proposition.
Further, we assume that $n=p^{k}$, $k\geq 2$. Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Note that $X_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $X_{D}=D^{\#}$ by Lemma 3.2(1). Since $X\neq X_{H}\times X_{D}$, we have $|X_{H}|>1$. So $|X_{H}|\in\{2,3\}$ by Lemma 7.8. Lemma 7.10 implies that $\lambda_{X}=1$. If $P\nleq\operatorname{rad}(X_{h,p^{l}})$ for some $h\in X_{H}$ and $l\in\{2,\ldots,k\}$, then the set $Y=X^{[p]}$, which is an $\mathcal{A}$-set by Lemma 2.4, lies outside $H\cup D$ and $Y_{D}\subsetneq D^{\#}$, a contradiction to Lemma 3.2(1). Therefore
@@
for every $h\in X_{H}$. Observe that
@@
because $X_{D}=D^{\#}$. Together with Lemma 2.5, this yields $\mathcal{B}=\operatorname{Cyc}(K,P)$, where $\mathcal{B}=\operatorname{Span}_{\mathbb{Z}}\{\underline{e},\underline{X}_{h,p}:~h\in X_{H}\}$ and $K$ is a subgroup of $\operatorname{Aut}(P)\cong C_{p-1}$ of index $|X_{H}|$, and hence
@@
If $|X_{H}|=2$ and (p1)/2(p-1)/2 is odd, then $\mathcal{B}$ is an antisymmetric $S$-ring of rank $3$. Using the formulas for the intersection numbers of an antisymmetric association scheme of rank $3$ (see, e.g. [3, Exercise 2.7.57]), one can compute that
@@
If $|X_{H}|=2$ and (p1)/2(p-1)/2 is even or $|X_{H}|=3$, then from [29, Lemma 5.2] it follows that

### v3

@@ 7.2. Nondense $S$-rings
    $4\nmid|G/L|$ and $|\operatorname{rad}(\mathcal{A}_{G/L})|=1$.
@@
Let $\widehat{\mathcal{A}}$ be the $S$-ring dual to $\mathcal{A}$ over $\widehat{G}\cong G$. Since $D$ is not an $\mathcal{A}$-subgroup, the group $D^{\bot}\cong E_{4}$ is not an $\widehat{\mathcal{A}}-$subgroup by Lemma 4.1(1). So Theorem 7.3 holds for $\widehat{\mathcal{A}}$. If $\widehat{\mathcal{A}}$ is trivial, then so is $\mathcal{A}$ and we are done. If $\widehat{\mathcal{A}}$ is a nontrivial tensor product, then so is $\mathcal{A}$ by Lemma 4.1(3) and we are done.
[…]
Lemma 4.1(2) implies that $\widehat{\mathcal{A}_{G/L}}=\widehat{\mathcal{A}}_{\widehat{U}}$ and $\widehat{\mathcal{A}_{U/L}}=\widehat{\mathcal{A}}_{\widehat{U}/\widehat{L}}$. So if $\widehat{L}$ is $\otimes$-complemented in $\widehat{\mathcal{A}}_{\widehat{U}}$, then $\mathcal{A}_{U/L}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$ by Lemma 4.1(3) as required.
[…]
7.3. Dense $S$-rings
The lemma below is straightforward.
@@ 7.3. Dense $S$-rings
Every $S$-ring over $H\cong E_{4}$ is normal and equals $\operatorname{Cyc}(K,H)$, where $K\leq\operatorname{Aut}(H)$ is of order at most $3$.
Remark 7.10.
In Lemma 7.9, we have $\operatorname{Cyc}(K_{H},H)=\mathbb{Z}H$ if $|K_{H}|=1$, $\operatorname{Cyc}(K_{H},H)\cong\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2}$ if $|K_{H}|=2$, and $\operatorname{Cyc}(K_{H},H)=\mathcal{T}_{H}$ if $|K_{H}|=3$.
Further throughout this subsection, we assume that $H$ and $D$ are $\mathcal{A}$-subgroups. Denote two distinct elements from $H^{\#}$ by $a$ and $b$. Given $X\in\mathcal{S}(\mathcal{A})$, put $\lambda_{X}=|X\cap Hx|$, where $x\in X$. Due to Lemma 2.1, $\lambda_{X}$ does not depend on $x$. Clearly, $1\leq\lambda_{X}\leq|X_{H}|\leq 3$.
Lemma 7.11.
Let $\mathcal{A}$ be dense and $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Then $X=X_{H}\times X_{D}$ or $\lambda_{X}=1$.
@@
Assume that $X\neq X_{H}\times X_{D}$. Then $\lambda_{X}<|X_{H}|$. We are done if $\lambda_{X}=1$. So it remains to consider only the case when $1<\lambda_{X}<|X_{H}|\leq 3$. Since $H\cong E_{4}$, this is possible only if $X_{H}=H^{\#}=\{a,b,ab\}$ and $\lambda_{X}=2$.
@@
Assume that two latter equalities hold. Then $Y=(X_{H}\times X_{D})\setminus X$ is an $\mathcal{A}$-set containing exactly one element from $Hx$ for every $x\in X_{D}$. Together with Lemma 3.2(1) and the equality $X_{H}=H^{\#}$, this implies that $Y$ is exactly a basic set of $\mathcal{A}$ with
@@
Let $x,x^{\prime}\in X_{D}$ be of the same order. Then there is a positive integer $m$ coprime to $2n$ such that $x^{m}=x^{\prime}$. Since $\lambda_{X}=2$, the sets $H_{x}=Xx^{-1}\cap H$ and $H_{x^{\prime}}=X(x^{\prime})^{-1}\cap H$ are subsets of $H^{\#}$ of size $2$. Therefore there is $h\in H_{x}\cap H_{x^{\prime}}$. One can see that $(hx)^{m}=hx^{\prime}\in X^{(m)}\cap X$. So $X^{(m)}=X$ by Lemma 2.3. This yields that
@@
Observe that $|Y_{h_{0}}|=|Y_{h}|$ by Lemma 2.1. In view of Eqs. (20) and (21), to obtain a contradiction, it suffices to show that
@@
The set $X_{D}$ is a basic set of $\mathcal{A}_{D}$ (Lemma 3.2(1)). If $|X_{D}\setminus D_{l}^{*}|=0$, then Eq. (22) is obvious. Suppose that $|X_{D}\setminus D_{l}^{*}|>0$, i.e. $X_{D}$ is nonregular. If $n=pq$, then Lemma 6.12 implies that $\mathcal{A}_{D}$ is trivial or a nontrivial generalized wreath product. So $X_{D}=D^{\#}$ or $X_{D}$ is a union of some $L$-cosets, where $L=\operatorname{rad}(X_{D})$ is a nontrivial $\mathcal{A}_{D}$-subgroup. In the former case, $|X_{D}\cap D_{l}^{*}|=(p-1)(q-1)$ and $|X_{D}\setminus D_{l}^{*}|=p+q-2$, whereas in the latter one $|X_{D}\cap D_{l}^{*}|=r(|L|-1)$ and $|X_{D}\setminus D_{l}^{*}|=r$, where $r=|X_{D}/L|$. If $n=p^{k}$, then by Lemma 6.11, we obtain $X_{D}=U\setminus L$ for some $\mathcal{A}$-subgroups $U\cong C_{p^{l}}>L$ and consequently $|X_{D}\cap D_{l}^{*}|=p^{l-1}(p-1)$ and $|X_{D}\setminus D_{l}^{*}|=p^{l-1}-|L|$. Since $p$ and $q$ are odd, Eq. (22) holds in all cases and we are done. ∎
Theorem 7.12.
Let $\mathcal{A}$ be dense. Suppose that $|\operatorname{rad}(\mathcal{A}_{D})|=1$ and $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$. Then:
@@
By the condition $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ and Lemma 3.2(2), we have $\mathcal{A}_{H}\neq\mathbb{Z}H$. Then Lemma 7.9 and Remark 7.10 imply that $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order $2$ or $3$, and a unique nonsingleton basic set of $\mathcal{A}_{H}$ is of size $|K_{H}|$. We divide the rest of the proof into two cases depending on whether $\mathcal{A}_{D}$ is cyclotomic or not.
Case 1: $\mathcal{A}_{D}$ is noncyclotomic. The condition $|\operatorname{rad}(\mathcal{A}_{D})|=1$ of the theorem and Lemma 6.2 imply that $\mathcal{A}_{D}$ is trivial whenever $n=p^{k}$. The same arguments and Lemma 6.12 imply that $\mathcal{A}_{D}$ is trivial in case $n=pq$. Thus, $\mathcal{A}_{D}$ is trivial in any case and hence $D^{\#}\in\mathcal{S}(\mathcal{A}_{D})$. Moreover, $n$ is not prime by Lemma 6.7. Therefore $n=pq$ or $n=p^{k}$ for $k\geq 2$.
Let $n=pq$. Then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$ and $D$ is the least $\mathcal{A}$-subgroup whose order is divisible by $p$. By Lemma 5.3 applied to $\mathcal{A}$, $H$, and $D$, we have $\mathcal{A}=\mathcal{A}_{H}\star\mathcal{A}_{D}$. Since $H\cap D$ is trivial, we obtain $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, a contradiction to the assumption of the theorem.
Now let $n=p^{k}$, $k\geq 2$. Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ by the condition of the theorem, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that
(23) $X\neq X_{H}\times X_{D}.$
Note that $X_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $X_{D}=D^{\#}$ by Lemma 3.2(1). Eq. (23) implies that $|X_{H}|>1$. So $|X_{H}|\in\{2,3\}$ by Lemma 7.9. Eq. (23) and Lemma 7.11 yield that $\lambda_{X}=1$. If $P\nleq\operatorname{rad}(X_{h,p^{l}})$ (see Notation for $X_{h,p^{l}}$) for some $h\in X_{H}$ and $l\in\{2,\ldots,k\}$, then the set $Y=X^{[p]}$, which is an $\mathcal{A}$-set by Lemma 2.4, lies outside $H\cup D$ and $Y_{D}\subsetneq D^{\#}$, a contradiction to Lemma 3.2(1). Therefore
@@
for every $h\in X_{H}$.
Lemma 2.5 applied to $X$, $h\in X_{H}$, and $l=p$ implies that the sets $X_{h,p}$, $h\in X_{H}$, are orbits of some $K\leq\operatorname{Aut}(P)$. This yields that all of these sets are of the same size $|K|$. Since $X_{D}=D^{\#}$, we have
@@
and hence
$\sum\limits_{h\in X_{H}}|X_{h,p}|=p-1.$
Thus,
@@
Further, we are going to compute $\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}$. Recall that $|X_{H}|\in\{2,3\}$. If $|X_{H}|=2$ and (p1)/2(p-1)/2 is odd, then $\operatorname{Cyc}(K,P)$ is an antisymmetric $S$-ring of rank $3$. Using the formulas for the intersection numbers of an antisymmetric association scheme of rank $3$ (see, e.g. [3, Exercise 2.7.57]), one can compute that
@@
If $|X_{H}|=2$ and (p1)/2(p-1)/2 is even or $|X_{H}|=3$, then $X_{h,p}=X_{h,p}^{-1}$ for every $h\in X_{H}$ and the required sum is a special case of the sum computed in [31, Lemma 5.2]. So applying this lemma, we obtain